OpenStudy (anonymous):

If the range of the function is f(x) = 5/x + 6.5 is {7, 6.83, 6.78, 6.75}, what is its domain? A) {5, 10, 15, 20} B) {10, 15, 18, 20} C) {10, 15, 17, 20} D) {10, 14, 16, 20} E) {8, 15, 18, 22} @agent0smith @nikato @Mertsj @julianassange

OpenStudy (mertsj):

They have given you y values. Those are f(x) values. Replace f(x) with each of the given range values and solve for x. Those x vales are the domain.

OpenStudy (mertsj):

I will illustrate with the first one.

OpenStudy (mertsj):

\[f(x)=\frac{5}{x}+6.5\] \[7=\frac{5}{x}+6.5\] \[.5=\frac{5}{x}\] \[.5x=5\] \[x=\frac{5}{.5}=10\]

OpenStudy (mertsj):

So one member of the domain is 10

OpenStudy (mertsj):

Find the others in the same way.

OpenStudy (anonymous):

Would the answer be B?

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

That's what I got.

OpenStudy (anonymous):

Ty (:

OpenStudy (anonymous):

thanks for putting your question on here cx it helps alot because i have the same one